# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

### v1

@@ Appendix B More examples
Single hypothesis; task: infer property; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-2 ontology tree.
@@
World Model: Rompuses are dalpists. Sarpers are dalpists. Each gergit is a dalpist. Pamela is a sarper. Jerry is a dalpist. Amy is a rompus.
Observations: Amy is rainy. Pamela is rainy. Jerry is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-3 ontology tree.
@@
World Model: Pamela is a porpor. Jerry is a dalpist. Twimpees are sarpers. Each gergit is a dalpist. Each storpist is a gergit. Amy is a storpist. Porpors are rompuses. Sarpers are dalpists. Rompuses are dalpists.
Observations: Jerry is not muffled. Amy is not muffled. Pamela is not muffled.
Hypotheses: Each dalpist is not muffled.
Single hypothesis; task: infer property; height-4 ontology tree.
@@
World Model: Amy is a felper. Jerry is a dalpist. Pamela is a felper. Each rompus is a dalpist. Each porpor is a rompus. Sarpers are dalpist s. Shumples are twimpees. Gergits are dalpists. Each frompor is a porpor. Storpists are gergits. Twimpees are sarpers. Felpers are storpists.
Observations: Amy is translucent. Pamela is translucent. Jerry is translucent.
Hypotheses: Each dalpist is translucent.
@@
Single hypothesis; task: infer membership relation; height-1 ontology tree.
World Model: Dalpists are liquid. Dalpists are rainy. Each dalpist is brown.
Observations: Jerry is brown. Jerry is rainy. Jerry is liquid.
Hypotheses: Jerry is a dalpist.
Single hypothesis; task: infer membership relation; height-2 ontology tree.
@@
World Model: Rompuses are dalpists. Jerry is a dalpist. Each sarper is a dalpist. Each dalpist is brown. Edward is a sarper. Gergits are dalpists. Gergits are rainy. Amy is a rompus. Dalpists are liquid.
Observations: Pamela is rainy. Pamela is liquid. Pamela is brown.
Hypotheses: Pamela is a gergit.
Single hypothesis; task: infer membership relation; height-3 ontology tree.
@@
World Model: Each sarper is a dalpist. Jerry is a dalpist. Barbara is a porpor. Each dalpist is moderate. Rompuses are dalpists. Twimpees are sarpers. Storpists are gergits. Each dalpist is large. Edward is a sarper. Pamela is a gergit. Gergits are dalpists. Debra is a storpist. Porpors are rompuses. Amy is a rompus. Twimpees are not muffled.
Observations: Nicole is moderate. Nicole is large. Nicole is not muffled.
Hypotheses: Nicole is a twimpee.
Single hypothesis; task: infer membership relation; height-4 ontology tree.
@@
World Model: Amy is a rompus. Each porpor is a rompus. Each sarper is a dalpist. Sharon is a felper. Rompuses are dalpists. Each shumple is a twimpee. Each dalpist is fruity. Barbara is a porpor. Every frompor is a porpor. Every gergit is a dalpist. Edward is a sarper. Raymond is a frompor. Each storpist is a gergit. Jerry is a dalpist. Each dalpist is moderate. Shumples are translucent. Felpers are storpists. Nicole is a twimpee. Pamela is a gergit. Twimpees are sarpers. Debra is a storpist.
Observations: Michael is fruity. Michael is translucent. Michael is moderate.
Hypotheses: Michael is a shumple.
@@
Single hypothesis; task: infer subtype relation; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is a rompus. Jerry is a rompus. Pamela is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-2 ontology tree.
@@
World Model: Gergits are dalpists. Porpors are dalpists. Each sarper is a dalpist. Pamela is a porpor. Jerry is a dalpist. Amy is a gergit.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-3 ontology tree.
@@
World Model: Amy is a storpist. Jerry is a dalpist. Frompors are porpors. Each sarper is a dalpist. Each twimpee is a sarper. Pamela is a twimpee. Storpists are gergits. Porpors are dalpists. Gergits are dalpists.
Observations: Jerry is a rompus. Amy is a rompus. Pamela is a rompus.
Hypotheses: Each dalpist is a rompus.
Single hypothesis; task: infer subtype relation; height-4 ontology tree.
@@
World Model: Each sarper is a dalpist. Jerry is a dalpist. Amy is a shumple. Pamela is a shumple. Each storpist is a gergit. Dumpuses are frompors. Gergits are dalpists. Felpers are storpists. Each twimpee is a sarper. Frompors are porpors. Shumples are twimpees. Porpors are dalpists.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Each dalpist is a rompus.
Multiple hypothesis; height-1 ontology tree.
@@
World Model: Debra is a dalpist. Amy is a dalpist. Pamela is a dalpist. Edward is a dalpist. Dalpists are liquid. Each dalpist is moderate. Each dalpist is brown. Barbara is a dalpist.
Observations: Barbara is rainy. Jerry is a rompus. Jerry is liquid. Jerry is brown. Pamela is a rompus. Edward is rainy. Debra is rainy. Jerry is moderate. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are rainy. Jerry is a dalpist.
Multiple hypothesis; height-2 ontology tree.
@@
World Model: Nicole is a gergit. Porpors are dalpists. Pamela is a sarper. Sharon is a gergit. Porpors are not cold. Dalpists are rainy. Raymond is a gergit. Each porpor is translucent. Each sarper is a dalpist. Jerry is a dalpist. Edward is a dalpist. Every gergit is a dalpist. Debra is a dalpist. Barbara is a sarper.
Observations: Barbara is liquid. Amy is translucent. Sharon is moderate. Debra is liquid. Amy is rainy. Raymond is moderate. Jerry is a rompus. Amy is not cold. Pamela is a rompus. Nicole is moderate. Edward is liquid. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are liquid. Every gergit is moderate. Amy is a porpor.
Multiple hypothesis; height-3 ontology tree.
@@
World Model: Barbara is a sarper. Sarpers are dalpists. Debra is a shumple. Every felper is a sarper. Raymond is a porpor. Sharon is a storpist. Each storpist is a gergit. Helen is a frompor. Porpors are dalpists. Each dalpist is liquid. Amy is a sarper. Each dalpist is salty. Felpers are rainy. Every shumple is a porpor. Each twimpee is a sarper. Nicole is a gergit. Each frompor is a sarper. Jerry is a dalpist. Dalpists are not muffled. Each gergit is a dalpist. Michael is a twimpee. Sarpers are amenable.
Observations: Debra is large. Amy is a rompus. Pamela is a rompus. Edward is liquid. Pamela is rainy. Edward is not muffled. Barbara is large. Pamela is not muffled. Edward is large. Edward is salty. Jerry is a rompus. Pamela is amenable.
Hypotheses: Each dalpist is a rompus. Dalpists are large. Edward is a dalpist. Pamela is a felper.
Multiple hypothesis; height-4 ontology tree.
@@
World Model: Every kurpor is a twimpee. Each gergit is pale. Patricia is a dumpus. Timples are frompors. Sharon is a folpee. George is an orgit. Rimpees are felpers. Stephen is a shumple. Mark is an orgit. Joseph is a porpor. Andrew is an orgit. Each tumpus is a felper. Gergits are dalpists. Felpers are not dark. Helen is a timple. Jason is a tumpus. Amy is a pergit. Rimpees are angry. Porpors are dalpists. Raymond is a twimpee. Each rimpee is moderate. Linda is a tumpus. Every shumple is a porpor. Each frompor is a sarper. Every dumpus is a storpist. Gergits are amenable. Barbara is an orgit. Rachel is a frompor. Felpers are sarpers. Jacob is a felper. Janet is a tumpus. Michael is a timple. Steven is a boompist. Edward is a dalpist. Storpists are gergits. Each boompist is a frompor. Debra is a dalpist. Nicole is a sarper. Pergits are felpers. Shirley is a timple. Jerry is a dalpist. Pame la is a kurpor. Every folpee is a frompor. Each sarper is a dalpist. Eric is a storpist. Orgits are shumples. Gergits are not o paque. Twimpees are sarpers.
Observations: Linda is brown. Helen is salty. George is translucent. Jason is brown. Raymond is large. Sharon is large. Jerry is a rompus. William is angry. Barbara is liquid. Edward is liquid. Emily is amenable. Janet is brown. Emily is not opaque. Shirley is salty. William is not dark. Pamela is a rompus. Nicole is large. William is moderate. Andrew is translucent. Michael is salty. Amy is a rompus. Debra is liquid. Emily is pale. Mark is translucent.
Hypotheses: Dalpists are rompuses. Every dalpist is liquid. Emily is a gergit. Sarpers are large. Timples are salty. William is a rimpee. Tumpuses are brown. Orgits are translucent
Appendix C Algorithms
Refer to Algorithm 1 and 2.
if task is infer property then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes
 generates the observations where selected members have the property
end if
if task is infer membership then
 select a random leaf node
 randomly select a property of the leaf node
 randomly select a property of the root node
 randomly select a property of the selected node’s non-root ancestors.
 generates the observations where this entity have all properties
end if
if task is infer subtype relation then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes

### v2

@@ Appendix C More examples of reasoning questions
Single hypothesis example; task: infer property; height-1 ontology tree. World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist. Observations: Amy is rainy. Jerry is rainy. Pamela is rainy. Hypotheses: Dalpists are rainy.
@@
Single hypothesis example; task: infer property; height-2 ontology tree. World Model: Rompuses are dalpists. Sarpers are dalpists. Each gergit is a dalpist. Pamela is a sarper. Jerry is a dalpist. Amy is a rompus. Observations: Amy is rainy. Pamela is rainy. Jerry is rainy. Hypotheses: Dalpists are rainy.
@@
Single hypothesis example; task: infer property; height-3 ontology tree. World Model: Pamela is a porpor. Jerry is a dalpist. Twimpees are sarpers. Each gergit is a dalpist. Each storpist is a gergit. Amy is a storpist. Porpors are rompuses. Sarpers are dalpists. Rompuses are dalpists. Observations: Jerry is not muffled. Amy is not muffled. Pamela is not muffled. Hypotheses: Each dalpist is not muffled.
@@
Single hypothesis example; task: infer property; height-4 ontology tree. World Model: Amy is a felper. Jerry is a dalpist. Pamela is a felper. Each rompus is a dalpist. Each porpor is a rompus. Sarpers are dalpists. Shumples are twimpees. Gergits are dalpists. Each frompor is a porpor. Storpists are gergits. Twimpees are sarpers. Felpers are storpists. Observations: Amy is translucent. Pamela is translucent. Jerry is translucent. Hypotheses: Each dalpist is translucent.
@@
Single hypothesis example; task: infer membership relation; height-1 ontology tree. World Model: Dalpists are liquid. Dalpists are rainy. Each dalpist is brown. Observations: Jerry is brown. Jerry is rainy. Jerry is liquid. Hypotheses: Jerry is a dalpist.
@@
Single hypothesis example; task: infer membership relation; height-2 ontology tree. World Model: Rompuses are dalpists. Jerry is a dalpist. Each sarper is a dalpist. Each dalpist is brown. Edward is a sarper. Gergits are dalpists. Gergits are rainy. Amy is a rompus. Dalpists are liquid. Observations: Pamela is rainy. Pamela is liquid. Pamela is brown. Hypotheses: Pamela is a gergit.
@@
Single hypothesis example; task: infer membership relation; height-3 ontology tree. World Model: Each sarper is a dalpist. Jerry is a dalpist. Barbara is a porpor. Each dalpist is moderate. Rompuses are dalpists. Twimpees are sarpers. Storpists are gergits. Each dalpist is large. Edward is a sarper. Pamela is a gergit. Gergits are dalpists. Debra is a storpist. Porpors are rompuses. Amy is a rompus. Twimpees are not muffled. Observations: Nicole is moderate. Nicole is large. Nicole is not muffled. Hypotheses: Nicole is a twimpee.
@@
Single hypothesis example; task: infer membership relation; height-4 ontology tree. World Model: Amy is a rompus. Each porpor is a rompus. Each sarper is a dalpist. Sharon is a felper. Rompuses are dalpists. Each shumple is a twimpee. Each dalpist is fruity. Barbara is a porpor. Every frompor is a porpor. Every gergit is a dalpist. Edward is a sarper. Raymond is a frompor. Each storpist is a gergit. Jerry is a dalpist. Each dalpist is moderate. Shumples are translucent. Felpers are storpists. Nicole is a twimpee. Pamela is a gergit. Twimpees are sarpers. Debra is a storpist. Observations: Michael is fruity. Michael is translucent. Michael is moderate. Hypotheses: Michael is a shumple.
@@
Single hypothesis example; task: infer subtype relation; height-1 ontology tree. World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist. Observations: Amy is a rompus. Jerry is a rompus. Pamela is a rompus. Hypotheses: Dalpists are rompuses.
@@
Single hypothesis example; task: infer subtype relation; height-2 ontology tree. World Model: Gergits are dalpists. Porpors are dalpists. Each sarper is a dalpist. Pamela is a porpor. Jerry is a dalpist. Amy is a gergit. Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus. Hypotheses: Dalpists are rompuses.
@@
Single hypothesis example; task: infer subtype relation; height-3 ontology tree. World Model: Amy is a storpist. Jerry is a dalpist. Frompors are porpors. Each sarper is a dalpist. Each twimpee is a sarper. Pamela is a twimpee. Storpists are gergits. Porpors are dalpists. Gergits are dalpists. Observations: Jerry is a rompus. Amy is a rompus. Pamela is a rompus. Hypotheses: Each dalpist is a rompus.
@@
Single hypothesis; task: infer subtype relation; height-4 ontology tree. World Model: Each sarper is a dalpist. Jerry is a dalpist. Amy is a shumple. Pamela is a shumple. Each storpist is a gergit. Dumpuses are frompors. Gergits are dalpists. Felpers are storpists. Each twimpee is a sarper. Frompors are porpors. Shumples are twimpees. Porpors are dalpists. Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus. Hypotheses: Each dalpist is a rompus.
@@
Multiple hypothesis example; height-1 ontology tree. World Model: Debra is a dalpist. Amy is a dalpist. Pamela is a dalpist. Edward is a dalpist. Dalpists are liquid. Each dalpist is moderate. Each dalpist is brown. Barbara is a dalpist. Observations: Barbara is rainy. Jerry is a rompus. Jerry is liquid. Jerry is brown. Pamela is a rompus. Edward is rainy. Debra is rainy. Jerry is moderate. Amy is a rompus. Hypotheses: Dalpists are rompuses. Dalpists are rainy. Jerry is a dalpist.
@@
Multiple hypothesis example; height-2 ontology tree. World Model: Nicole is a gergit. Porpors are dalpists. Pamela is a sarper. Sharon is a gergit. Porpors are not cold. Dalpists are rainy. Raymond is a gergit. Each porpor is translucent. Each sarper is a dalpist. Jerry is a dalpist. Edward is a dalpist. Every gergit is a dalpist. Debra is a dalpist. Barbara is a sarper. Observations: Barbara is liquid. Amy is translucent. Sharon is moderate. Debra is liquid. Amy is rainy. Raymond is moderate. Jerry is a rompus. Amy is not cold. Pamela is a rompus. Nicole is moderate. Edward is liquid. Amy is a rompus. Hypotheses: Dalpists are rompuses. Dalpists are liquid. Every gergit is moderate. Amy is a porpor.
@@
Multiple hypothesis example; height-3 ontology tree. World Model: Barbara is a sarper. Sarpers are dalpists. Debra is a shumple. Every felper is a sarper. Raymond is a porpor. Sharon is a storpist. Each storpist is a gergit. Helen is a frompor. Porpors are dalpists. Each dalpist is liquid. Amy is a sarper. Each dalpist is salty. Felpers are rainy. Every shumple is a porpor. Each twimpee is a sarper. Nicole is a gergit. Each frompor is a sarper. Jerry is a dalpist. Dalpists are not muffled. Each gergit is a dalpist. Michael is a twimpee. Sarpers are amenable. Observations: Debra is large. Amy is a rompus. Pamela is a rompus. Edward is liquid. Pamela is rainy. Edward is not muffled. Barbara is large. Pamela is not muffled. Edward is large. Edward is salty. Jerry is a rompus. Pamela is amenable. Hypotheses: Each dalpist is a rompus. Dalpists are large. Edward is a dalpist. Pamela is a felper.
@@
Multiple hypothesis example; height-4 ontology tree. World Model: Every kurpor is a twimpee. Each gergit is pale. Patricia is a dumpus. Timples are frompors. Sharon is a folpee. George is an orgit. Rimpees are felpers. Stephen is a shumple. Mark is an orgit. Joseph is a porpor. Andrew is an orgit. Each tumpus is a felper. Gergits are dalpists. Felpers are not dark. Helen is a timple. Jason is a tumpus. Amy is a pergit. Rimpees are angry. Porpors are dalpists. Raymond is a twimpee. Each rimpee is moderate. Linda is a tumpus. Every shumple is a porpor. Each frompor is a sarper. Every dumpus is a storpist. Gergits are amenable. Barbara is an orgit. Rachel is a frompor. Felpers are sarpers. Jacob is a felper. Janet is a tumpus. Michael is a timple. Steven is a boompist. Edward is a dalpist. Storpists are gergits. Each boompist is a frompor. Debra is a dalpist. Nicole is a sarper. Pergits are felpers. Shirley is a timple. Jerry is a dalpist. Pamela is a kurpor. Every folpee is a frompor. Each sarper is a dalpist. Eric is a storpist. Orgits are shumples. Gergits are not opaque. Twimpees are sarpers. Observations: Linda is brown. Helen is salty. George is translucent. Jason is brown. Raymond is large. Sharon is large. Jerry is a rompus. William is angry. Barbara is liquid. Edward is liquid. Emily is amenable. Janet is brown. Emily is not opaque. Shirley is salty. William is not dark. Pamela is a rompus. Nicole is large. William is moderate. Andrew is translucent. Michael is salty. Amy is a rompus. Debra is liquid. Emily is pale. Mark is translucent. Hypotheses: Dalpists are rompuses. Every dalpist is liquid. Emily is a gergit. Sarpers are large. Timples are salty. William is a rimpee. Tumpuses are brown. Orgits are translucent.
Appendix D Details of GPT-4o paraphrase
D.1 Prompts
We present our prompts for GPT-4o to paraphrase synthetic reasoning questions into more natural ones in Figure 6.
Prompt for Question Rewriting Instruction We prompt GPT-4o to rewrite synthetic reasoning questions into more natural, human-readable language while preserving the original semantics and reasoning structure. Prompt Template You are a professional editor specializing in rewriting synthetic reasoning questions into natural, human-readable language. Rewrite the following question to be clear, fluent, and natural while preserving all logical structure and meaning. Do not simplify the reasoning steps, remove constraints, or change the problem’s difficulty. Return only the improved question with the exact same semantics, but in more natural language. Input question: {{QUESTION}}
Figure 6: Prompt used to paraphrase synthetic reasoning questions into more natural language.
D.2 Manual inspection
We also manually verify that the generated questions from GPT-4o are sound by randomly sampling 200 reasoning questions of varying difficulty and manually checking that GPT-4o’s output does not alter the problem’s semantics. We conduct this study and iteratively improve the prompt until achieving 100% accuracy.
Appendix E Details of hypothesis quality metric
E.1 Theoretical justification
Our quality metric can be motivated by an argument from Bayesian decision theory. Let $T$ be a theory which consists of a set of observations $o_{1},\ldots,o_{n}$ and a set of proofs $\pi_{1},\ldots,\pi_{n}$ where each proof $\pi_{i}$ explains the observation $o_{i}$ (i.e., the conclusion of $\pi_{i}$ is $o_{i}$). Let
$p(T)=\prod_{i=1}^{n}p(\pi_{i}\mid\pi_{1},\ldots,\pi_{i-1})\,p(o_{i}\mid\pi_{i})$
be a prior distribution over the theory. Ideally, during inductive/abductive reasoning, given a new observation $o_{n+1}$, we aim to find the proof $\pi_{n+1}$ that maximizes the posterior
$p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n},o)$ $\propto p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n})$
$\quad\cdot p(o_{n+1}\mid\pi_{n+1})$
Since
$p(o_{i}\mid\pi_{i})=\mathds{1}\{o_{i}\text{ is the conclusion of }\pi_{i}\},$
the posterior is equal to
$p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n})$
whenever the conclusion of $\pi_{n+1}$ is $o_{n+1}$.
Now suppose we make the mild assumption that $p(T)$ is such that $p(T)>p(T^{\prime})$ whenever $T$ has fewer axioms in its proofs than does $T^{\prime}$. Therefore, $p(T)$ is maximized if $\pi_{n+1}$ minimally introduces new axioms (and equivalently, maximally reuses existing axioms). Note that the quality metric $g(\pi_{n+1})$ is the ratio of two quantities: (1) the average number of times each premise $h\in\pi_{n+1}$ appears in $\pi_{1},\ldots,\pi_{n}$, and (2) the average number of times each premise in the ground-truth proof appears in $\pi_{1},\ldots,\pi_{n}$. We specifically selected each ground-truth proof to be the unique proof with the minimum number of new axioms. If we consider two proofs $\pi_{n+1}$ and $\pi^{\prime}_{n+1}$ such that $\pi_{n+1}$ has more new axioms than $\pi^{\prime}_{n+1}$, then the sum in the numerator of $g(\pi_{n+1})$ would contain more zero terms than that of $g(\pi^{\prime}_{n+1})$, and so
$g(\pi_{n+1})<g(\pi^{\prime}_{n+1})$
necessarily.
Thus, with mild assumptions on the prior over the theory $p(T)$, our proposed quality metric provides an easy-to-compute proxy for the posterior probability of the theory, where more probable theories will have higher quality scores. Furthermore, our quality metric is agnostic to the specific choice of the prior $p(T)$.
E.2 Human validation experiments
To evaluate whether the proposed metric aligns with human judgments of hypothesis quality, we conduct a human evaluation study. We recruit three independent annotators with graduate-level STEM degrees in US R1 universities. The evaluation dataset contains 100 reasoning questions, each associated with three candidate hypotheses. For each question, annotators are asked to select the hypothesis that better explains observations and is most concise. The presentation order of hypotheses is randomized to avoid positional bias. Each reasoning question is evaluated by three annotators, and the final preference label is determined by majority vote.
To quantify alignment with human judgments, we measure agreement accuracy, defined as the percentage when the proposed metric assigns the highest score to the human-preferred hypothesis. The proposed metric achieves an agreement of 79%, substantially outperforming a random baseline (33.3%) and a length-based heuristic baseline (51%). Human annotations exhibit substantial agreement (Fleiss’ $\kappa$ = 0.75), indicating that the evaluation task is reasonably well-defined and that the proposed metric captures meaningful aspects of hypothesis quality.
Please refer to Figure 7 for the instruction to annotators.
Instruction for annotators --- You will be given a set of reasoning questions. Each question includes: * A description of observations * Three candidate hypotheses that attempt to explain these observations --- Your task is to select the single best hypothesis according to the following criteria: * The hypothesis should account for the given observations coherently and logically. * The hypothesis should be as concise as possible. --- Important Guidelines * Select only one hypothesis per question. * Do not choose based on writing style, fluency, or formatting alone. * Focus on the validity and quality of hypotheses, not superficial features. * The order of hypotheses is randomized; treat each option independently. * If multiple hypotheses seem similar, choose the one that is most concise. --- Your selection will be used to determine the hypothesis that best explains the observations.
Figure 7: Instruction for annotators.
Appendix F Algorithms
Algorithms 1 and 2 describe how we generate ontology trees and hypotheses for the examples in our framework.
1function generate_hypothesis(ontology tree $T$)
2    if task is “infer property”
3       randomly select a member of the root node of $T$
4       randomly select a member of a leaf node of $T$
5       generate the observations where selected members have the property
6
7   else if task is “infer membership”
8       select a random leaf node of $T$
9       randomly select a property of the leaf node
10       randomly select a property of the root node
11       randomly select a property of the selected node’s non-root ancestors
12       generate observations where this entity has all properties
13
14   else if task is “infer subtype relation”
15       randomly select a member of the root node of $T$
16       randomly select a member of a leaf node of $T$
17       randomly select a member of a non-leaf of $T$